Sales (measured by number of boxes sold with 12 balls in each box) of six different...

  1. Sales (measured by number of boxes sold with 12 balls in each box) of six different brands of golf balls at a retail store were recorded for one month. The observed (absolute) frequencies and expected (absolute) frequencies as well as the MegaStat results of a χ2(Chi-square) Goodness of Fit test appear below.

observed expected

Brand 1: Bridgestone

131

200

Brand 2: Callaway

273

200

Brand 3: Nike

119

200

Brand 4: Srixon

301

200

Brand 5: Titleist

176

200

Brand 6: Wilson

200

200

total

1200

1200

137.14

chi-square

5

df

7.25E-28

p-value

  1. (3 pts) Write out the Ho for this Goodness of Fit test.
  1. (4 pts.) Interpret the results (for α = .05). Can we reject Ho? Explain.
Answer #1

(a)

Ho: The proportion of boxes sold are equal for all 6 brands. i.e. 1/6

Pbridgestone = 1/6, Pcallaway = 1/6, Pnike = 1/6, Psrixon = 1/6, Ptitlesit = 1/6, Pwilson = 1/6

Ha: Some of the population proportions differ from the values stated in the null hypothesis

This corresponds to a Chi-Square test for Goodness of Fit.

(b)

\chi^{2} = 137.14

\chi^{2} critical = 11.070

As \chi^{2} (137.14) is greater than \chi^{2} critical, we reject the Null hypothesis.

Also as the p value ( 7.25E-28) is less than \alpha (0.05) we reject the Null hypothesis.

Hence we have sufficient evidence to believe some of the population proportions differ from the values stated in the null hypothesis.